OpenStudy (anonymous):

Paul throws a baseball onto the field. The baseball leaves Paul’s hand with an initial upward velocity of 15 meters per second. If the baseball is 1 meter off the ground when it leaves Paul’s hand, about how long will it take for the baseball to hit the ground? Round your answer to the nearest tenth.

OpenStudy (anonymous):

s = -1, u = 15, t = ?, a = -9.8 m/s/s s = ut + 0.5 at^2 -1 = 15t + 4.9t^2 4.9t^2 + 15t + 1 = 0 solve for t

OpenStudy (anonymous):

sorry - typo \ equation should be 4.9t^2 - 15t - 1 = 0

OpenStudy (anonymous):

t = 3.1 secs to nearest tenth second

OpenStudy (anonymous):

r u there cherry?

OpenStudy (anonymous):

hello!!